# Leptogenesis with perturbations in type-II and type-III seesaw models

### v2

@@ I Introduction
and therefore one can speculate that the correction will tend to be larger (due to the well established mathematical inequality $2>1$). This is relevant for standard DM self-annihilations during freeze-out, and also for type-II Ma and Sarkar (1998); Hambye and Senjanovic (2004); Antusch and King (2004); Hambye et al. (2006) and type-III leptogenesis Hambye et al. (2004); Strumia (2009); Hambye (2012). In this paper, we study the latter two leptogenesis models as examples of acoustically driven freeze-out involving annihilations, and leave DM freeze-out for another day Baldes (2025) (annihilations can also play an important role in simple extensions of the type-I scenario Aristizabal Sierra et al. (2014), but we only consider the minimal scenario here).
@@
One objection to this whole enterprise may be that without a reason for expecting significant enhancements in $\mathcal{R}_{i}$ — at length scales of order of the hubble horizon $\sim 1/H$ during the freeze-out process — these calculations are only of academic interest, and will be unlikely to be of relevance for explaining the observed universe, unlike the perturbations at CMB scales Senatore et al. (2009) (for historical studies of small scale inhomogeneties during BBN, see e.g. Terasawa and Sato (1989); Kurki-Suonio and Matzner (1989)). We note, however, that strong phase transitions associated with the generation of heavy DM Hambye and Strumia (2013); Hambye et al. (2018); Baldes and Garcia-Cely (2019); Baldes et al. (2021a, 2022); Kierkla et al. (2023); Wong and Xie (2023); Balan et al. (2025) or heavy (seesaw) states associated with baryogensis Lazarides et al. (1986); Buchmuller et al. (2012); Buchmüller et al. (2013); Brivio and Trott (2017, 2019); Brdar et al. (2019a, b, c); Brivio et al. (2019); Brdar et al. (2019d); Baldes et al. (2021b); Huang and Xie (2022); Dasgupta et al. (2022); Chun et al. (2023), will also lead to the generation of curvature perturbations with length scales $\sim 1/H$ Liu et al. (2023); Elor et al. (2024); Lewicki et al. (2024); Zou et al. (2025). If the reheating temperature is subsequently high enough for the DM or heavy seesaw states to come back into equilibrium, the generated perturbations may well then lead to acoustically driven freeze-out playing a role.
@@ II Leptogenesis models
$\mathcal{L}\supset-\frac{1}{2}(y_{{TL}}L^{T}Ci\tau_{2}TL+y_{{TH}}^{\dagger}M_{T}\tilde{H}^{T}i\tau_{2}T\tilde{H}+H.c.),$ (4)
@@
where $B_{L}+B_{H}=1$ are the branching fractions. For simplicity, when we scan over parameter space, solving the Boltzmann equations below, we will be considering the limit $B_{L}=B_{H}=0.5$. Then
@@ IV.1 Type-I Boltzmann equation
with $g_{\ast}=106.75$ effective degrees-of-freedom. The baryon-minus lepton asymmetry normalized to entropy is denoted $Y_{\mathcal{B}-\mathcal{L}}$. The equilibrium number density of heavy neutrinos in the Maxwell-Boltzmann approximation is given by
@@ IV.2 Type-II Boltzmann equation
Figure 3: Contours showing the enhancement of the yield with perturbations for the three seesaw models (with $B_{L}=B_{H}=0.5$ for type-II). As well established in the previous literature, preferred regions (green) assuming hierarchical spectrum CP violation limits, Eqs. (18)-(20), are at somewhat higher $M_{1}$ in the type-II and type-III scenarios, due to the gauge interactions keeping the number densities closer to equilibrium. (We are assuming initial thermal $N$ abundance for the type-I green region, which affects the shape of this region at small $\tilde{m}_{1}$.) The new result shows that in the presence of perturbations, i.e. acoustically driven freeze-out, gauge interactions allow yield enhancements in the weak washout regime, $\tilde{m_{1}}<m_{\ast}$. This occurs in parameter ranges where the decay time approximately matches the freeze-out time of the gauge interactions. At large values of $M_{1}$ the yields become suppressed due to off-shell $2\to 2$ washout interactions. For very large $M_{1}$ the couplings become non-perturbative (red shaded regions).
@@ V Results
Finally in Fig. 3, we collect our results, by showing standard plots of the preferred regions for leptogenesis in the different seesaw models, together with contours showing the enhancement of the yield for acoustically driven freeze-out, in terms of the parameters $M_{1}$ and $\tilde{m_{1}}$. Cosmological observations disfavour $\sum m_{\nu i}>0.12\;\mathrm{eV}$ Aghanim et al. (2020); Alam et al. (2021), so also $\tilde{m}_{1}>0.12\;\mathrm{eV}$ is disfavoured Buchmuller et al. (2003), but we extend the plots to higher values for completeness. At large values of $M_{1}$ the couplings required to explain the neutrino masses become non-perturbative. Observations of the CMB limit the post-inflationary reheat temperature to $T_{\rm RH}<6.6\times 10^{15}\;\mathrm{GeV}$ Akrami et al. (2020); Ade et al. (2021) which gives a similar consistency constraint on the scenario, $M_{1}<T_{\rm RH}$, as perturbativity.
@@
In the strong washout regime, $\tilde{m_{1}}>m_{\ast}$, departure from equilibrium is set by the decay rate, making the three scenarios qualitatively similar. In the weak washout regime, $\tilde{m_{1}}<m_{\ast}$, freeze-out of the gauge annihilations plays the central role in determining the departure from equilibrium in type-II and type-III, leading to significant qualitative differences between these scenarios and type-I leptogenesis. This was already known to be the case for the behaviour of the evolution of the particle number densities and asymmetries without the density perturbations. The effects of the acoustically driven freeze-out follows a similar pattern.
@@ VI Conclusion
We have extended the study of acoustically driven freeze out to include $2\to 2$ interactions, by examining in some detail leptogenesis in the type-II and type-III seesaw models in the presence of perturbations, and contrasting our results with those of type-I leptogenesis. Modest enhancements of the yield, $\sim 40\%$, are possible with primordial curvature perturbations $\mathcal{R}_{i}\sim 0.3$, already in the weak washout regime in the scenarios with gauge interactions. (The crucial ingredient being annihilations involving the species freezeing out in the external states — similar effects could be expected in extensions of the type-I seesaw involving annihilations of $N_{1}$ into scalars Aristizabal Sierra et al. (2014).)
Assuming the existence of such primordial small scale perturbations, the size of the effect is of the same order as the effect of the Sommerfeld enhancment in type-II and III leptogenesis Strumia (2009), although most relevant for somewhat differing areas of parameter space.
It would be of interest to study acoustically driven freeze-out in the context of dark matter, to see what effect, if any, is present in the freeze-out of stable species. Preliminary investigations suggest that in the standard DM freeze-out scenario, modest enhancements of the required DM annihilation cross section $\sim 10\%$ are necessary to match the relic abundance, for perturbations $\mathcal{R}_{i}\sim 0.3$ Baldes (2025) (for effects of small scale perturbations during freeze-in, see (Stebbins, 2023, Fig. 1)). It would also be of interest to see whether models featuring inflation, or strong early universe phase transitions, can be found which link physics of the particles undergoing freeze-out, with the generation of large $\mathcal{R}_{i}$ on the length scales $\sim 1/H(\bar{T}=M)$, relevant for acoustically driven freeze-out.

### v3

@@ I Introduction
and therefore one can speculate that the correction will tend to be larger (due to the well established mathematical inequality $2>1$). This is relevant for standard DM self-annihilations during freeze-out, and also for type-II Ma and Sarkar (1998); Hambye and Senjanovic (2004); Antusch and King (2004); Hambye et al. (2006) and type-III leptogenesis Hambye et al. (2004); Strumia (2009); Hambye (2012). In this paper, we study the latter two leptogenesis models as examples of acoustically driven freeze-out involving annihilations, and leave DM freeze-out for another day Baldes and Jinno (2025) (annihilations can also play an important role in simple extensions of the type-I scenario Aristizabal Sierra et al. (2014), but we only consider the minimal scenario here).
@@
One objection to this whole enterprise may be that without a reason for expecting significant enhancements in $\mathcal{R}_{i}$ — at length scales of order of the hubble horizon $\sim 1/H$ during the freeze-out process — these calculations are only of academic interest, and will be unlikely to be of relevance for explaining the observed universe, unlike the perturbations at CMB scales Senatore et al. (2009) (for historical studies of small scale inhomogeneties during BBN, see e.g. Terasawa and Sato (1989); Kurki-Suonio and Matzner (1989), and for more recent studies of the effect of density perturbations on BBN see Jeong et al. (2014); Nakama et al. (2014); Inomata et al. (2016)). We note, however, that strong phase transitions associated with the generation of heavy DM Hambye and Strumia (2013); Hambye et al. (2018); Baldes and Garcia-Cely (2019); Baldes et al. (2021a, 2022); Kierkla et al. (2023); Wong and Xie (2023); Balan et al. (2025) or heavy (seesaw) states associated with baryogensis Lazarides et al. (1986); Buchmuller et al. (2012); Buchmüller et al. (2013); Brivio and Trott (2017, 2019); Brdar et al. (2019a, b, c); Brivio et al. (2019); Brdar et al. (2019d); Baldes et al. (2021b); Huang and Xie (2022); Dasgupta et al. (2022); Chun et al. (2023), will also lead to the generation of curvature perturbations with length scales $\sim 1/H$ Liu et al. (2023); Elor et al. (2024); Lewicki et al. (2024); Zou et al. (2025); Franciolini et al. (2025). If the reheating temperature is subsequently high enough for the DM or heavy seesaw states to come back into equilibrium, the generated perturbations may well then lead to acoustically driven freeze-out playing a role.
@@ II Leptogenesis models
$\mathcal{L}\supset-\frac{1}{2}(y_{{TL}}L^{T}Ci\tau_{2}TL+y_{{TH}}M_{T}\tilde{H}^{T}i\tau_{2}T\tilde{H}+\mathrm{H.c.}),$ (4)
@@
where $B_{L}+B_{H}=1$ are the branching fractions. One thus has
$\Gamma_{T}=\frac{\tilde{m_{1}}M_{1}^{2}}{16\pi v^{2}\sqrt{B_{L}B_{H}}},\quad\mathrm{[Type-II}].$ (13)
For simplicity, when we scan over parameter space, solving the Boltzmann equations below, we will be considering the limit $B_{L}=B_{H}=0.5$. Then
@@ IV.1 Type-I Boltzmann equation
with $g_{\ast}=106.75$ effective degrees-of-freedom. The baryon-minus-lepton asymmetry normalized to entropy is denoted $Y_{\mathcal{B}-\mathcal{L}}$. The equilibrium number density of heavy neutrinos in the Maxwell-Boltzmann approximation is given by
@@ IV.2 Type-II Boltzmann equation
Figure 3: Contours showing the enhancement of the yield with perturbations for the three seesaw models (with $B_{L}=B_{H}=0.5$ for type-II). The new result shows that in the presence of perturbations, i.e. acoustically driven freeze-out, gauge interactions allow yield enhancements in the weak washout regime, $\tilde{m_{1}}<m_{\ast}$. This occurs in parameter ranges where the decay time approximately matches the freeze-out time of the gauge interactions. As well established in the previous literature, preferred regions (green) assuming hierarchical spectrum CP violation limits, Eqs. (19)-(21), are at somewhat higher $M_{1}$ in the type-II and type-III scenarios, due to the gauge interactions keeping the number densities closer to equilibrium. The green regions are shown without the perturbations. The increase in the yield, due to the perturbations, will only lead to small changes at the borders of these areas as can be inferred by the blue contours. (The assumption of the initial $N$ abundance for the type-I seesaw affects the preferred region at small $\tilde{m}_{1}$.) At large values of $M_{1}$ the yields become suppressed due to off-shell $2\to 2$ washout interactions. For very large $M_{1}$ the couplings become non-perturbative (red shaded regions).
@@ V Results
Finally in Fig. 3, we collect our results, by showing standard plots of the preferred regions for leptogenesis in the different seesaw models (without the perturbations), together with contours showing the enhancement of the yield for acoustically driven freeze-out, in terms of the parameters $M_{1}$ and $\tilde{m_{1}}$. Cosmological observations disfavour $\sum m_{\nu i}>0.12\;\mathrm{eV}$ Aghanim et al. (2020); Alam et al. (2021), so also $\tilde{m}_{1}>0.12\;\mathrm{eV}$ is disfavoured Buchmuller et al. (2003), but we extend the plots to higher values for completeness. At large values of $M_{1}$ the couplings required to explain the neutrino masses become non-perturbative. Observations of the CMB limit the post-inflationary reheat temperature to $T_{\rm RH}<6.6\times 10^{15}\;\mathrm{GeV}$ Akrami et al. (2020); Ade et al. (2021) which gives a similar consistency constraint on the scenario, $M_{1}<T_{\rm RH}$, as perturbativity.
As initial conditions we typically assume thermal abundance of the heavy see-saw states and zero initial asymmetry. For the type-I seesaw, the final asymmetry is sensitive to both these initial conditions in the weak washout regime. We show in Fig. 3 the effect of assuming a highly suppressed abundance, $Y_{N}(\bar{z}=0.01)=10^{-12}\,Y_{N}^{\rm eq}$, on the preferred region for hierarchical spectra (relevant at small $\tilde{m_{1}}$). The sensitivity is outside the parameter region in which the acoustically driven freeze-out plays a role. For type-II and type-III, as the gauge interactions bring the heavy states to their thermal abundance values, there is no dependence on the assumed initial abundance of the heavy state (only depedence on the assumed zero initial asymmetry in the weak washout regime).
@@
In the strong washout regime, $\tilde{m_{1}}>m_{\ast}$, departure from equilibrium is set by the decay rate, making the three scenarios qualitatively similar. In the weak washout regime, $\tilde{m_{1}}<m_{\ast}$, freeze-out of the gauge annihilations plays the central role in determining the departure from equilibrium in type-II and type-III, leading to significant qualitative differences between these scenarios and type-I leptogenesis. This was already known to be the case for the behaviour of the evolution of the particle number densities and asymmetries without the density perturbations. The effects of the acoustically driven freeze-out follow a similar pattern.
@@ VI Conclusion
We have extended the study of acoustically driven freeze out to include $2\to 2$ interactions, by examining in some detail leptogenesis in the type-II and type-III seesaw models in the presence of perturbations, and contrasting our results with those of type-I leptogenesis. Modest enhancements of the yield, $\sim 40\%$, are possible with primordial curvature perturbations $\mathcal{R}_{i}\sim 0.3$, already in the weak washout regime in the scenarios with gauge interactions. (The crucial ingredient being annihilations involving the species freezing out in the external states — similar effects could be expected in extensions of the type-I seesaw involving annihilations of $N_{1}$ into scalars Aristizabal Sierra et al. (2014).)
Assuming the existence of such primordial small scale perturbations, the size of the effect is of the same order as the effect of the Sommerfeld enhancment in type-II and III leptogenesis Strumia (2009) (which we have also included in our calculation), although most relevant for somewhat differing areas of parameter space.
Other sizable corrections we have not included are the following: If the heavy states have small mass splittings, resonance effects can lead to significant increases in the CP violation, and hence lower the allowed regions to the $\sim$ TeV scale in both the type-I Pilaftsis and Underwood (2004) and type-II/III cases Strumia (2009). Flavour effects can also be significant, improving the efficiency by around an order of magnitude, and thus lowering the minimum seesaw scale for hierarchical leptogenesis in all seesaw cases also by around an order of magnitude Abada et al. (2006a); Nardi et al. (2006b); Abada et al. (2006b); Lavignac and Schmauch (2015). Spectator processes, not included here, which are in thermal equilibrium during leptogenesis lead to $\mathcal{O}(1)$ differences in the yield Buchmuller and Plumacher (2001); Nardi et al. (2006a).
For the range of effective neutrino masses considered here, indeed for $\tilde{m}_{1}\gtrsim 10^{-7.5}$ eV, the heavy neutrinos do not come to dominate the universe before their decay (even in the type-I case assuming an initial thermal abundance), so there is no early epoch of matter domination. Once well inside the Horizon and well after asymmetry generation Jeong et al. (2014); Nakama et al. (2014), however, diffusion damping of the oscillations will lead to entropy production, at second order in the perturbations Chluba et al. (2012a, b). This increases the entropy density and thus leads to a suppression of the yield by a factor $s_{f}/s_{i}=(1+3\langle\delta_{T}^{2}\rangle$). For our monochromatic perturbations, the time and $\delta$ phase average is $\langle\delta_{T}^{2}\rangle=|\mathcal{R}_{i}|^{2}/4$, so the correction is $<7\%$ for the perturbations $|\mathcal{R}_{i}|\leq 0.3$ considered here. (Larger effects are possible for broad, scale invariant enhancements of the perturbation spectra at small scales Jeong et al. (2014); Nakama et al. (2014); Inomata et al. (2016).)
It would be of interest to study acoustically driven freeze-out in the context of dark matter, to see what effect, if any, is present in the freeze-out of stable species. Preliminary investigations suggest that in the standard DM freeze-out scenario, modest enhancements of the required DM annihilation cross section $\sim 10\%$ are necessary to match the relic abundance, for perturbations $\mathcal{R}_{i}\sim 0.3$ Baldes and Jinno (2025) (for effects of small scale perturbations during freeze-in, see (Stebbins, 2023, Fig. 1)). It would also be of interest to see whether models featuring inflation, or strong early universe phase transitions, can be found which link physics of the particles undergoing freeze-out, with the generation of large $\mathcal{R}_{i}$ on the length scales $\sim 1/H(\bar{T}=M)$, relevant for acoustically driven freeze-out.